OpenStudy (anonymous):

Not understanding this!!! Simplify: y^2 × y^-4

OpenStudy (johnweldon1993):

It's important to remember that when multiplying exponents with the same base...you leave the bases alone and ADD the exponents...

OpenStudy (anonymous):

Ohh... I just don't understand it

OpenStudy (johnweldon1993):

Alright...so here....the "base" would be 'y'... \[\large y^2 \times y^{-4} = y^{2 + (-4)} = y^{2 - 4} = y^{-2}\] Does that make sense?

OpenStudy (anonymous):

Aha so the answer would be Y^2?

OpenStudy (johnweldon1993):

y^-2 ...did I make that -2 too small up there? sorry about that....but yeah it would be y^(2 - 4) so that is y^-2

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (johnweldon1993):

No problem!